OpenStudy (hy123):

Let (5,-4) be a point on the terminal side of (theta). Find the exact values of cos, csc, and tan.

OpenStudy (hy123):

Hey gane

OpenStudy (hy123):

Site seems like it is running smooth now?

OpenStudy (baldymcgee6):

(x,y) ---> (5,-4) r=sqrt(5^2-(-4)^2)= sqrt(9)=3 cos(x)=x/r cox(x)=5/3 x= 1.098612288668109691395245236922525704647490557822749 csc(x)= 1/sin(x) sin(x)=y/r sin(x)=-4/3 csc(x)=1/(-4/3) tan(x) = y/x tan(x) = -4/5 x=-0.67474094222355266305652097360981361507400625484071